OpenStudy (anonymous):

free medals for answers!!!!!!!!!!!!!!!!!!!!!! The sum of Anjie's age and Carol's age is 66. Carol's age is 2 more than 3 times Anjie's age. How old are Carol and Anjie???????

OpenStudy (anonymous):

Please show work,,,, FREE MEDALS!!!!!!! you have to actually answer the question tho :)

OpenStudy (anonymous):

medals are free anyhow

OpenStudy (anonymous):

Carol is 45 years old.

OpenStudy (anonymous):

Solve the following for a and c. {a + c == 66, c == 2 a + 3}

OpenStudy (anonymous):

how old is Anjie?

OpenStudy (anonymous):

a + c = 66 a + 45 = 66 a = 66 - 45 a = 66 - 45 a = 21 Anjie is 21 years old.